Question Help
















A bus traveled on a level road for 
5 
hours at an average speed 20 miles per hour faster than it traveled on a winding road. The time spent on the winding road was 
3 
hours. Find the average speed on the level road if the entire trip was 
364 
miles.















A bus traveled on a level road for 
5 
hours at an average speed 20 miles per hour faster than it traveled on a winding road. The time spent on the winding road was 
3 
hours. Find the average speed on the level road if the entire trip was 
364 
miles.

Answer by ankor@dixie-net.com(22740)

What's the point of using up so much space?
:
A bus traveled on a level road for 5hours at an average speed 20 miles per hour faster than it traveled on a winding road.
The time spent on the winding road was 3 hours.
Find the average speed on the level road if the entire trip was 364 miles.
:
let s = the speed on the winding road
then
(s+20) = the speed on a level road
:
Write a distance equation; dist = time * speed
level dist + winding dist = 364 mi
5(s+20) + 3s = 364
5s + 100 + 3s = 364
5s + 3s = 364 - 100
8s = 264
s = 264/8
s = 33 mph on the winding road
:
I'll let you find the speed on the level road, check your solutions in the 1st equation